GIF: ATL-CMS medium gas system
From July 19th :
-
test regulation box with LOGO Patrick + Student w28:Present status/results
-
O2 analysis LPh (if I understood, we have three and the order for the closed loops was delayed due to some issue/mistake at cern purchase office level )
-
mixer injection to be moved before humidifier (it allows operation with humid gas in open mode) Fred to check implication of modification Status/update
-
modification to check calibration of O2 sensor LPh Status/update
-
Check if old RPC MFCs R&D gas can be upgraded to 3 times the flow Patrick Status/update
-
Purifier:
-
Time needed for elec modifications (0.5 weks) and chassie (1.5 weeks) total 10 days expected by end w33 Patrick
-
Update plan of the area in order to evaluate cable length Roberto
-
Mechanics (0.5 weeks) expected by w32 Fred
-
Order FESTO to replace spares used for modif purify Fred to follow with Andrea
-
Software commissioning w34 Steven
-
Medium Closed loop rack:
-
Mechanically expected to be completed by w30 Fred Staus/Update
-
Electrical part 1.5 weeks for the rack only (Purif detailed above) Expected by wXYZ (I will update on Monday) Patrick Status/Update
-
Mixer & Humidifier:
-
Mechanically expected to be completed by w30 Fred Staus/Update
-
Electrical part 1 weeks for the rack only (Purif detailed above) Expected by wXYZ (I will update on Monday) Patrick Status/Update
-
MFCs should be back on w35 Patrick Status/Update
-
Did we received the MFCs for Humidifier? Patrick Status/Update
-
Installation and commissioning
-
Expected by wXYZ (I will update on Monday) Fred
From June 23rd :
-
test regulation box with LOGO Patrick + Student w28
-
check displays (how many needed and how many available?) Beatrice confirmed all ok for all 3 systems
-
H2O analysers confirmed all ok for all systems (Beatrice)
-
O2 analysis Beatrice to check with LPh
-
mixer injection to be moved before humidifier (it allows operation with humid gas in open mode) Fred to check implication of modification
-
modification to check calibration of O2 sensor LPh
-
Check if old RPC MFCs R&D gas can be upgraded to 3 times the flow Patrick
From June 6th
-
2 MFCs for ATL_CMS_RPC R134a 300 nl/h low DP to buy Patrick
-
VCR connectors for 6 gas mixers & humidifier Albin
-
O2, H2O analysers for cosmic trigger closed loop have been ordered? Beatrice
-
P&ID purifier ex TGC. Find drawing à Fred
-
How to operate old exTGC purify at GIF. Meeting Patrick, Steven, Fred, Roberto
-
Output 4-20 mA and GSM LOGO to be ordered Patrick
-
Check PI regulation LOGO for mixer 6 gas R&D Patrick(or Steven)
Hello,
I propose to review the small closed loop and mixers status, plan, …. All what is needed.
Is Thursday June 23th 2pm in the lab (face to face with the racks) fine with you?
Thanks a lot
Roberto
Hello,
Update status, excel tables orders, …
In order of priority:
-
Plan item 59 (expected to be completed by w13)
Mixer R&D GAS SYSTEM (6 GASES) [ATEX] – BC 33175
HUMIDIFIER R&D GAS SYSTEM [NO ATEX] – BC 33321
-
Plan item 60 (expected to be completed by w16)
Mixer GEM (BO) [NO ATEX] – BC 33321
-
Plan item 60 (expected to be completed by w16)
Mixer RPC (BO) [ATEX] – BC 33321
HUMIDIFIER RPC (BO) [NO ATEX] – BC 33321
-
Plan item 60 (expected to be completed by w16)
Mixer ATL-CMS-RPC-GIF [ATEX] – BC 33321
HUMIDIFIER ATL-CMS-RPC-GIF [NO ATEX] – BC 33321
-
Plan item 61 (expected to be completed by w29)
SCLOOP R&D GAS SYSTEM [NO ATEX] – BC 33321
-
Plan item 61 (expected to be completed by w29)
MEDIUM CLOOP ATL-CMS-RPC-GIF [NO ATEX] – BC 33321
-
Plan item 61 (expected to be completed by w29)
SCLOOP COSMIC TRIGGER GIF [NO ATEX] – BC 33321
To be done:
Update excel table for orders à Beatrice, Patrick, Fred
Update drawing closed-loops (3 drawings) à Fred
Let me know if I forgot something.
Thanks a lot
Roberto
Hello,
I propose to meet Monday 9am to review the folders for the small closed loop
Let me know if it is fine with you. Please bring all the info/drawings you already have.
Thanks a lot
Roberto